# Computing a Fixed Point of Contraction Maps in Polynomial Queries

### v1

@@ 5 Existence of Balanced Point
so intuitively $x$ is far from $\mathcal{P}_{i^{*}}(p^{*},+1)$. From this we can conclude that $B(x,1/t_{\ell})\cap\mathcal{P}_{i^{*}}(p^{t_{\ell}},+1)$ is empty. To see this is the case, for any $y\in B(x,1/t_{\ell})$, it follows from $\|x-y\|_{\infty}\leq\|x-y\|_{2}\leq 0.1$ and $\|p^{*}-p^{t_{\ell}}\|_{\infty}\leq 0.1$ and (6) that
$y_{i^{*}}-y_{j}\leq p^{t_{\ell}}_{i^{*}}-p^{t_{\ell}}_{j}-(1-0.4)<p^{t_{\ell}}_{i^{*}}-p^{t_{\ell}}_{j}$
and thus, $y\notin\mathcal{P}_{i^{*}}(p^{t_{\ell}},+1)$. The other case follows from a similar argument.
@@
Now we consider the general case when $A$ is not empty and let $\alpha>0$ be the smallest value in $A$; note that $\alpha\leq 1/2$. In this case we let $\ell$ be a sufficiently large integer such that $1/t_{\ell}\leq 0.1\alpha$ and $\|p^{t_{\ell}}-p^{*}\|_{\infty}\leq 0.1\alpha$. Similarly it suffices to show that every point $x\in T$ satisfies that
$\mathcal{P}_{i^{*}}(p^{t_{\ell}},+1)\cap B(x,1/t_{\ell})\neq\emptyset\ \ \Longrightarrow\ \ x\in\mathcal{P}_{i^{*}}(p^{*},+1).$
@@
Let’s prove the contrapositive so take any $x\notin\mathcal{P}_{i^{*}}(p^{*},+1)$. There exists a $j\neq i^{*}$ such that either
$x_{i^{*}}-p^{*}_{i^{*}}<x_{j}-p^{*}_{j}\quad\text{or}\quad x_{i^{*}}-p^{*}_{i^{*}}<p^{*}_{j}-x_{j}$
@@
For the first case we have $x_{i^{*}}-x_{j}<p^{*}_{i^{*}}-p^{*}_{j}$. Since $p^{*}_{i^{*}}-p^{*}_{j}$ is either an integer or an integer $\pm$ something that is between $\alpha$ and $1-\alpha$, we have
$x_{i^{*}}-x_{j}\leq p^{*}_{i^{*}}-p^{*}_{j}-\alpha.$
@@
Given $p^{*}$, we round it to an integer point $q^{*}\in\texttt{OE}(n,k)$ as follows. First let $q_{i^{*}}^{*}\in[0,n]$ be an integer such that $|p^{*}_{i^{*}}-q^{*}_{i^{*}}|\leq 1/2$ (note that $q^{*}_{i^{*}}$ may not be unique but we can break ties arbitrarily). It is clear that $q_{i^{*}}^{*}\in\left\{0,\ldots,n\right\}$. Assume without loss of generality that $q_{i^{*}}^{*}$ is even (so we need to set $q_{j}^{*}$ to be even for every other $j$, in order to have $q^{*}\in\texttt{OE}(n,k)$). Then for each $j\neq i^{*}$:
  1. 1.
    $q^{*}_{j}=0$ if $p^{*}_{j}\in[-1/{4},0)$;
  2. 2.
    $q^{*}_{j}$ is set to be the even number in $\left\{n-1,n\right\}$ if $p^{*}_{j}\in(n,n+{1}/{4}]$; and
  3. 3.
    Otherwise, set $q_{j}^{*}$ to be an even number in $\{0,\ldots,n\}$ such that $|p^{*}_{j}-q_{j}^{*}|\leq 1$ (again breaking ties arbitrarily).
Note that we have $q^{*}\in\texttt{OE}(n,k)$ and it satisfies $|p^{*}_{j}-q_{j}^{*}|\leq{5}/{4}$ for all $j\neq i^{*}$.
@@
The point $q^{*}$ satisfies $q^{*}\in\emph{{OE}}(n,k)$ and
$\mathcal{P}_{i^{*}}(p^{*},+1)\cap T\subseteq\mathcal{P}_{i^{*}}(q^{*},+1)\cap T\quad\text{and}\quad\mathcal{P}_{i^{*}}(p^{*},-1)\cap T\subseteq\mathcal{P}_{i^{*}}(q^{*},-1)\cap T.$
@@
Let’s prove the first part since the other part is symmetric.
Let $x\in\texttt{OE}(n,k)$ be a point in $\mathcal{P}_{i^{*}}(p^{*},+1)\cap T$. So for every $j\neq i^{*}$ we have
$x_{i^{*}}-p^{*}_{i^{*}}\geq x_{j}-p^{*}_{j}\quad\text{and}\quad x_{i^{*}}-p^{*}_{i^{*}}\geq p^{*}_{j}-x_{j},$
@@
$x_{i^{*}}-x_{j}\geq p^{*}_{i^{*}}-p^{*}_{j}\quad\text{and}\quad x_{i^{*}}+x_{j}\geq p^{*}_{i^{*}}+p^{*}_{j}.$
@@
$x_{i^{*}}-x_{j}\geq q^{*}_{i^{*}}-q^{*}_{j}\quad\text{and}\quad x_{i^{*}}+x_{j}\geq q^{*}_{j}+q^{*}_{i^{*}}$
To see the first part, we have
$x_{i^{*}}-x_{j}\geq p^{*}_{i^{*}}-p^{*}_{j}\geq(q^{*}_{i^{*}}-1/2)-(q^{*}_{j}+5/4)=(q^{*}_{i^{*}}-q^{*}_{j})-7/4.$
But given that $x\in\texttt{OE}(n,k)$ and $q^{*}\in\texttt{OE}(n,k)$, both $x_{i^{*}}-x_{j}$ and $q^{*}_{i^{*}}-q^{*}_{j}$ are even numbers and thus, the inequality above implies $x_{i^{*}}-x_{j}\geq q^{*}_{i^{*}}-q^{*}_{j}.$ The other part can be proved similarly. ∎
@@ 6 Impossibility of Strong Approximation under Non-expansion
In the proof it will be more convenient to use as the domain a square that is tilted by $45^{\circ}$. We call a rectangle whose sides are at $45^{\circ}$ and $-45^{\circ}$, a diamond. Let $D$ be the diamond whose vertices are the midpoints of the sides of the unit square. Any function $g$ over $D$ can be extended to a function $g^{\prime}$ over the unit square, by defining for every point $p\in[0,1]^{2}$ the value of the function as $g^{\prime}(p)=g(\pi(p))$, where $\pi(p)$ is the projection of $p$ onto $D$. Clearly, for any two points $p,q\in[0,1]^{2}$, $\|\pi(p)-\pi(q)\|_{\infty}\leq\|p-q\|_{\infty}$, hence if the function $g$ over $D$ is non-expansive, then so is the function $g^{\prime}$ over $[0,1]^{2}$. Furthermore, the fixed points of $g^{\prime}$ are exactly the fixed points of $g$.
@@
We will show now that $f_{\delta,s}$ is a non-expansive function, i.e. that $\|f_{\delta,s}(p)-f_{\delta,s}(q)\|_{\infty}\leq\|p-q\|_{\infty}$ for all $p,q\in D$. We show first that it suffices to check pairs $p,q$ that are diagonal to each other, i.e. such that the line connecting them is at $45^{\circ}$ or $-45^{\circ}$. Note that such points have the property that the $L_{\infty}$ distance is tight in both coordinates, $\|p-q\|_{\infty}=|p_{1}-q_{1}|=|p_{2}-q_{2}|$.
@@
Let $x,y$ be any two points that are not diagonal. Consider the diamond with opposite vertices $x,y$, i.e. draw the lines through $x,y$ at $45^{\circ}$ and $-45^{\circ}$ and considered the rectangle enclosed by them. Let $z,w$ be the other two vertices of this diamond. Suppose without loss of generality that $\|x-y\|_{\infty}=x_{1}-y_{1}>|x_{2}-y_{2}|$. Then $x_{1}>z_{1}>y_{1}$, and similarly for $w$. We have $\|x-y\|_{\infty}=x_{1}-y_{1}=(x_{1}-z_{1})+(z_{1}-y_{1})$ $=\|x-z\|_{\infty}+\|z-y\|_{\infty}$. Since $f$ is non-expansive on diagonal pairs, $\|f(x)-f(z)\|_{\infty}\leq\|x-z\|_{\infty}$ and $\|f(z)-f(y)\|_{\infty}\leq\|z-y\|_{\infty}$. Therefore, $\|f(x)-f(y)\|_{\infty}\leq\|f(x)-f(z)\|_{\infty}+\|f(z)-f(y)\|_{\infty}$ $\leq\|x-z\|_{\infty}+\|z-y\|_{\infty}=\|x-y\|_{\infty}$. ∎
@@
So suppose $p,q$ are both in $D_{0}$. Assume first that the line $pq$ has angle $45^{\circ}$, i.e. $pq$ is parallel to the line $l_{0}$. Then $\|p-q\|_{\infty}=|pq|/\sqrt{2}$. Let $p^{\prime},q^{\prime}$ be the projections of $p,q$ on $l_{0}$, and let $p"=f_{\delta,s}(p)$, $q"=f_{\delta,s}(q)$. Then $|p^{\prime}p"|=(\delta-|pp^{\prime}|)\cdot|p^{\prime}s|$, $|q^{\prime}q"|=(\delta-|qq^{\prime}|)\cdot|q^{\prime}s|$. Since $pq$ is parallel to $l_{0}$, $(\delta-|pp^{\prime}|)=(\delta-|qq^{\prime}|)$ and $|pq|=|p^{\prime}q^{\prime}|$. Assume without loss of generality that $|p^{\prime}s|>|q^{\prime}s|$. Then $|p"q"|=|p^{\prime}q^{\prime}|-(\delta-|pp^{\prime}|)(|p^{\prime}s|-|q^{\prime}s|)\leq|p^{\prime}q^{\prime}|=|pq|$. Since $\|p-q\|_{\infty}=|pq|/\sqrt{2}$ and $\|f(p)-f(q)\|_{\infty}=|p"q"|/\sqrt{2}$, it follows that $\|f(p)-f(q)\|_{\infty}\leq\|p-q\|_{\infty}$.
Assume now that the line $pq$ has angle $-45^{\circ}$, i.e., $pq$ is perpendicular to $l_{0}$. Again $\|p-q\|_{\infty}=|pq|/\sqrt{2}$. Now $p$ and $q$ have the same projection $p^{\prime}=q^{\prime}$ on $l_{0}$. Let $p"=f_{\delta,s}(p)$, $q"=f_{\delta,s}(q)$. We have $|p^{\prime}p"|=(\delta-|pp^{\prime}|)\cdot|p^{\prime}s|$, and $|q^{\prime}q"|=|p^{\prime}q"|=(\delta-|qp^{\prime}|)\cdot|p^{\prime}s|$. Therefore, $|p"q"|=|(|pp^{\prime}|-|qp^{\prime}|)|\cdot|p^{\prime}s|$. If $p,q$ are on the same side of $l_{0}$ then $|(|pp^{\prime}|-|qp^{\prime}|)|=|pq|$. If $p,q$ are on opposite sides of $l_{0}$ then $|(|pp^{\prime}|-|qp^{\prime}|)|<|pq|$. In either case, we have $|p"q"|\leq|pq|\cdot|p^{\prime}s|<|pq|$, since $|p^{\prime}s|\leq|st|=1/\sqrt{2}$. Again, since $\|p-q\|_{\infty}=|pq|/\sqrt{2}$ and $\|f(p)-f(q)\|_{\infty}=|p"q"|/\sqrt{2}$, it follows that $\|f(p)-f(q)\|_{\infty}\leq\|p-q\|_{\infty}$. ∎
@@
Suppose there is a (randomized) algorithm $B$ that computes a point that is within 1/4 of a fixed point of a non-expansive function $f$ over the domain $D$ within a finite expected number $n$ of queries (the expectation is over the random choices of the algorithm). We will show how to solve faster the array search problem. Partition the diamond $D$ into $N=2^{2n}$ strips by drawing $N-1$ parallel lines at $45^{\circ}$, spaced at distance $1/(N\sqrt{2})$ from each other, between the NW and SE side of $D$. Let $S_{1},\ldots,S_{N}$ be the $N$ strips. Fix a $\delta<1/(N2\sqrt{2})$. For each $x\in[N]$, let $s_{x}$ be the point on the SW side of $S_{x}$ at Euclidean distance $\delta$ from the S vertex of $S_{x}$, and $t_{x}$ the point on the NE side of $S_{x}$ at Euclidean distance $\delta$ from the N vertex. Note that $\|s_{x}-t_{x}\|_{\infty}>1/2$. Let ${\mathcal{F}}$ be the family of non-expansive functions $\{f_{\delta,s_{x}},f_{\delta,t_{x}}~{}|x\in[N]\}$.

### v2

@@ 5 Existence of Balanced Point
so intuitively $x$ is far from $\mathcal{P}_{i}(p^{*},+1)$. From this we can conclude that $B(x,1/t_{\ell})\cap\mathcal{P}_{i}(p^{t_{\ell}},+1)$ is empty. To see this is the case, for any $y\in B(x,1/t_{\ell})$, it follows from $\|x-y\|_{\infty}\leq\|x-y\|_{2}\leq 0.1$ and $\|p^{*}-p^{t_{\ell}}\|_{\infty}\leq 0.1$ and (4) that
$y_{i}-y_{j}\leq p^{t_{\ell}}_{i}-p^{t_{\ell}}_{j}-(1-0.4)<p^{t_{\ell}}_{i}-p^{t_{\ell}}_{j}$
and thus, $y\notin\mathcal{P}_{i}(p^{t_{\ell}},+1)$.
The other case where $x_{i}-p^{*}_{i}<p^{*}_{j}-x_{j}$ is similar. ∎
Notice that $B(x,1/t_{\ell})$ are all disjoint from each other, so for every $x\in T$, we have
$\frac{\textbf{vol}\left(\left(\bigcup_{i\in[k]}\mathcal{P}_{i}(p^{t_{\ell}},s_{i})\right)\cap B(x,1/t_{\ell})\right)}{\textbf{vol}(S^{t_{\ell}})}\leq\frac{\textbf{vol}(B(x,1/t_{\ell}))}{\textbf{vol}(S^{t_{\ell}})}=\frac{1}{|T|}.$
Thus, Claim 1 implies that
$\frac{\left|\left(\bigcup_{i\in[k]}\mathcal{P}_{i}(p^{*},s_{i})\right)\cap T\right|}{|T|}\geq\frac{\textbf{vol}\left(\left(\bigcup_{i\in[k]}\mathcal{P}_{i}(p^{t_{\ell}},s_{i})\right)\cap S^{t_{\ell}}\right)}{\textbf{vol}(S^{t_{\ell}})}.$ (5)
It only remains to note that
$\textbf{vol}\left(\left(\bigcup_{i\in[k]}\mathcal{P}_{i}(p^{t_{\ell}},s_{i})\right)\cap S^{t_{\ell}}\right)=\sum_{i\in[k]}\textbf{vol}\left(\mathcal{P}_{i}(p^{t_{\ell}},s_{i})\cap S^{t_{\ell}}\right)$
and
$\textbf{vol}(\mathcal{P}_{i}(p^{t_{\ell}},+1)\cap S^{t_{\ell}})=\textbf{vol}(\mathcal{P}_{i}(p^{t_{\ell}},-1)\cap S^{t_{\ell}})$
for all $i\in[k]$. The first equation holds because the intersection of any two different pyramids $\mathcal{P}_{i}(p^{t_{\ell}},s_{i})$ lies on a hyperplane and thus has 0 volume, and the second equation follows from the choice of $p^{t_{\ell}}$. So Equation 5 implies Lemma 7 since
$\frac{\left|\left(\bigcup_{i\in[k]}\mathcal{P}_{i}(p^{*},s_{i})\right)\cap T\right|}{|T|}\geq\frac{\textbf{vol}\left(\left(\bigcup_{i\in[k]}\mathcal{P}_{i}(p^{t_{\ell}},s_{i})\right)\cap S^{t_{\ell}}\right)}{\textbf{vol}(S^{t_{\ell}})}=\sum_{i\in[k]}\frac{\textbf{vol}(\mathcal{P}_{i}(p^{t_{\ell}},s_{i})\cap S^{t_{\ell}})}{\textbf{vol}(S^{t_{\ell}})}=\frac{1}{2}.$
@@
Now we consider the general case when $A$ is not empty and let $\alpha>0$ be the smallest value in $A$; note that $\alpha\leq 1/2$. In this case we let $\ell$ be a sufficiently large integer such that $1/t_{\ell}\leq 0.1\alpha$ and $\|p^{t_{\ell}}-p^{*}\|_{\infty}\leq 0.1\alpha$. Similarly it suffices to show that for every point $x\in T$ and $i\in[k]$, the following property holds
$\mathcal{P}_{i}(p^{t_{\ell}},s_{i})\cap B(x,1/t_{\ell})\neq\emptyset\ \ \Longrightarrow\ \ x\in\mathcal{P}_{i}(p^{*},s_{i}).$
@@
Let’s prove the contrapositive so take any $x\in T$ and $i\in[k]$ such that $x\notin\mathcal{P}_{i}(p^{*},s_{i})$. Assume without loss of generality that $s_{i}=+1$. There exists a $j\neq i$ such that either
$x_{i}-p^{*}_{i}<x_{j}-p^{*}_{j}\quad\text{or}\quad x_{i}-p^{*}_{i}<p^{*}_{j}-x_{j}.$
@@
For the first case we have $x_{i}-x_{j}<p^{*}_{i}-p^{*}_{j}$. Since $p^{*}_{i}-p^{*}_{j}$ is either an integer or an integer $\pm$ something that is between $\alpha$ and $1-\alpha$, we have
$x_{i}-x_{j}\leq p^{*}_{i}-p^{*}_{j}-\alpha.$
@@
Given $p^{*}$, we round it to an integer point $q^{*}\in[0:n]^{k}$ as follows: Let $q_{i}^{*}\in[0,n]$ be an integer such that $|p^{*}_{i}-q^{*}_{i}|\leq 1/2$ for every $i\in[k]$ (in the case that $q^{*}_{i}$ is not unique, we break ties by choosing $q^{*}_{i}\in\left\{p^{*}_{i}-1/2,p^{*}_{i}+1/2\right\}$ such that $q^{*}_{i}$ is odd). It is clear that $q^{*}\in[0:n]^{k}$.
@@
The point $q^{*}$ satisfies $q^{*}\in[0:n]^{k}$ and
$\mathcal{P}_{i}(p^{*},+1)\cap T\subseteq\mathcal{P}_{i}(q^{*},+1)\cap T\quad\text{and}\quad\mathcal{P}_{i}(p^{*},-1)\cap T\subseteq\mathcal{P}_{i}(q^{*},-1)\cap T\quad\text{for all }i\in[k].$
@@
Fix an arbitrary $i\in[k]$. We show $\mathcal{P}_{i}(p^{*},+1)\cap T\subseteq\mathcal{P}_{i}(q^{*},+1)\cap T$ below, and the other part $\mathcal{P}_{i}(p^{*},-1)\cap T\subseteq\mathcal{P}_{i}(q^{*},-1)\cap T$ can be handled in the same way.
Let $x\in\texttt{EVEN}(n,k)$ be a point in $\mathcal{P}_{i}(p^{*},+1)\cap T$. So for every $j\neq i$ we have
$x_{i}-p^{*}_{i}\geq x_{j}-p^{*}_{j}\quad\text{and}\quad x_{i}-p^{*}_{i}\geq p^{*}_{j}-x_{j},$
@@
$x_{i}-x_{j}\geq p^{*}_{i}-p^{*}_{j}\quad\text{and}\quad x_{i}+x_{j}\geq p^{*}_{i}+p^{*}_{j}.$
@@
$x_{i}-x_{j}\geq q^{*}_{i}-q^{*}_{j}\quad\text{and}\quad x_{i}+x_{j}\geq q^{*}_{i}+q^{*}_{j}.$
Note that
$p_{i}^{*}\geq q^{*}_{i}-1/2\quad\text{and}\quad p_{j}^{*}\leq q_{j}^{*}+1/2.$
So we have
$x_{i}-x_{j}\geq p^{*}_{i}-p^{*}_{j}\geq(q^{*}_{i}-1/2)-(q^{*}_{j}+1/2)=(q^{*}_{i}-q^{*}_{j})-1.$
Given the rounding rule for obtaining $q^{*}$ and the fact that $x\in\texttt{EVEN}(n,k)$, we will argue that it must be the case where $x_{i}-x_{j}\geq q^{*}_{i}-q^{*}_{j}.$ Consider two cases below.
If $p_{i}^{*}>q^{*}_{i}-1/2$ or $p_{j}^{*}<q_{j}^{*}+1/2.$, then we know that $x_{i}-x_{j}>q^{*}_{i}-q^{*}_{j}-1$. Since both $x_{i}-x_{j}$ and $q^{*}_{i}-q^{*}_{j}$ are integer numbers, this implies $x_{i}-x_{j}\geq q^{*}_{i}-q^{*}_{j}.$
If $p_{i}^{*}=q^{*}_{i}-1/2$ and $p_{j}^{*}=q_{j}^{*}+1/2,$ then we know that both $q_{i}^{*}$ and $q_{j}^{*}$ are odd numbers, which means the parity of $|q_{i}^{*}-q_{j}^{*}|$ is even. Since the parity of $|x_{i}-x_{j}|$ is also even, $x_{i}-x_{j}\geq q^{*}_{i}-q^{*}_{j}-1$ implies $x_{i}-x_{j}\geq q^{*}_{i}-q^{*}_{j}$.
The other part $x_{i}+x_{j}\geq q^{*}_{i}+q^{*}_{j}$ can be proved similarly. ∎
@@ 6 Impossibility of Strong Approximation under Non-expansion
In the proof it will be more convenient to use as the domain a square that is tilted by $45^{\circ}$. We call a rectangle whose sides are at $45^{\circ}$ and $45^{\circ}$, a diamond. Let $D$ be the diamond whose vertices are the midpoints of the sides of the unit square. Any function $g$ over $D$ can be extended to a function $g^{\prime}$ over the unit square, by defining for every point $p\in[0,1]^{2}$ the value of the function as $g^{\prime}(p)=g(\pi(p))$, where $\pi(p)$ is the projection of $p$ onto $D$, namely, the point obtained by translating $p$ along lines perpendicular to the sides of $D$ until it intersects the boundary of the diamond. Clearly, for any two points $p,q\in[0,1]^{2}$, $\|\pi(p)-\pi(q)\|_{\infty}\leq\|p-q\|_{\infty}$, hence if the function $g$ over $D$ is non-expansive, then so is the function $g^{\prime}$ over $0,1]^{2}$. Furthermore, the fixed points of $g^{\prime}$ are exactly the fixed points of $g$.
@@
We will show now that $f_{\delta,s}$ is a non-expansive function, i.e. that $\|f_{\delta,s}(p)-f_{\delta,s}(q)\|_{\infty}\leq\|p-q\|_{\infty}$ for all $p,q\in D$. We show first that it suffices to check pairs $p,q$ that are diagonal to each other, i.e. such that the line connecting them is at $45^{\circ}$ or $45^{\circ}$. Note that such points have the property that the $L_{\infty}$ distance is tight in both coordinates, $\|p-q\|_{\infty}=|p_{1}-q_{1}|=|p_{2}-q_{2}|$.
@@
Let $x,y$ be any two points that are not diagonal. Consider the diamond with opposite vertices $x,y$, i.e. draw the lines through $x,y$ at $45^{\circ}$ and $45^{\circ}$ and consider the rectangle enclosed by them. Let $z,w$ be the other two vertices of this diamond. Suppose without loss of generality that $\|x-y\|_{\infty}=x_{1}-y_{1}>|x_{2}-y_{2}|$. Then $x_{1}>z_{1}>y_{1}$, and similarly for $w$. We have $\|x-y\|_{\infty}=x_{1}-y_{1}=(x_{1}-z_{1})+(z_{1}-y_{1})$ $\|x-z\|_{\infty}+\|z-y\|_{\infty}$. Since $f$ is non-expansive on diagonal pairs, $\|f(x)-f(z)\|_{\infty}\leq\|x-z\|_{\infty}$ and $\|f(z)-f(y)\|_{\infty}\leq\|z-y\|_{\infty}$. Therefore, $\|f(x)-f(y)\|_{\infty}\leq\|f(x)-f(z)\|_{\infty}+\|f(z)-f(y)\|_{\infty}$ $\leq\|x-z\|_{\infty}+\|z-y\|_{\infty}=\|x-y\|_{\infty}$. ∎
@@
So suppose $p,q$ are both in $D_{0}$. Assume first that the line $pq$ has angle $45^{\circ}$, i.e. $pq$ is parallel to the line $l_{0}$. Then $\|p-q\|_{\infty}=|pq|/\sqrt{2}$. Let $p^{\prime},q^{\prime}$ be the projections of $p,q$ on $l_{0}$, and let $p^{\prime\prime}=f_{\delta,s}(p)$, $q^{\prime\prime}=f_{\delta,s}(q)$. Then $|p^{\prime}p^{\prime\prime}|=(\delta-|pp^{\prime}|)\cdot|p^{\prime}s|$, $|q^{\prime}q^{\prime\prime}|=(\delta-|qq^{\prime}|)\cdot|q^{\prime}s|$. Since $pq$ is parallel to $l_{0}$, $(\delta-|pp^{\prime}|)=(\delta-|qq^{\prime}|)$ and $|pq|=|p^{\prime}q^{\prime}|$. Assume without loss of generality that $|p^{\prime}s|>|q^{\prime}s|$. Then $|p^{\prime\prime}q^{\prime\prime}|=|p^{\prime}q^{\prime}|-(\delta-|pp^{\prime}|)(|p^{\prime}s|-|q^{\prime}s|)\leq|p^{\prime}q^{\prime}|=|pq|$. Since $\|p-q\|_{\infty}=|pq|/\sqrt{2}$ and $\|f(p)-f(q)\|_{\infty}=|p^{\prime\prime}q^{\prime\prime}|/\sqrt{2}$, it follows that $\|f(p)-f(q)\|_{\infty}\leq\|p-q\|_{\infty}$.
Assume now that the line $pq$ has angle $45^{\circ}$, i.e., $pq$ is perpendicular to $l_{0}$. Again $\|p-q\|_{\infty}=|pq|/\sqrt{2}$. Now $p$ and $q$ have the same projection $p^{\prime}=q^{\prime}$ on $l_{0}$. Let $p^{\prime\prime}=f_{\delta,s}(p)$, $q^{\prime\prime}=f_{\delta,s}(q)$. We have $|p^{\prime}p^{\prime\prime}|=(\delta-|pp^{\prime}|)\cdot|p^{\prime}s|$, and $|q^{\prime}q^{\prime\prime}|=|p^{\prime}q^{\prime\prime}|=(\delta-|qp^{\prime}|)\cdot|p^{\prime}s|$. Therefore, $|p^{\prime\prime}q^{\prime\prime}|=|(|pp^{\prime}|-|qp^{\prime}|)|\cdot|p^{\prime}s|$. If $p,q$ are on the same side of $l_{0}$ then $|(|pp^{\prime}|-|qp^{\prime}|)|=|pq|$. If $p,q$ are on opposite sides of $l_{0}$ then $|(|pp^{\prime}|-|qp^{\prime}|)|<|pq|$. In either case, we have $|p^{\prime\prime}q^{\prime\prime}|\leq|pq|\cdot|p^{\prime}s|<|pq|$, since $|p^{\prime}s|\leq|st|=1/\sqrt{2}$. Again, since $\|p-q\|_{\infty}=|pq|/\sqrt{2}$ and $\|f(p)-f(q)\|_{\infty}=|p^{\prime\prime}q^{\prime\prime}|/\sqrt{2}$, it follows that $\|f(p)-f(q)\|_{\infty}\leq\|p-q\|_{\infty}$. ∎
@@
Suppose there is a (randomized) algorithm $B$ that computes a point that is within 1/4 of a fixed point of a non-expansive function $f$ over the domain $D$ within a finite expected number $n$ of queries (the expectation is over the random choices of the algorithm). We will show how to solve faster the array search problem. Partition the diamond $D$ into $N=2^{2n}$ strips by drawing N1N-1 parallel lines at $45^{\circ}$, spaced at distance $1/(N\sqrt{2})$ from each other, between the NW and SE side of $D$. Let $S_{1},\ldots,S_{N}$ be the $N$ strips. Fix a $\delta<1/(N2\sqrt{2})$. For each $x\in[N]$, let $s_{x}$ be the point on the SW side of $S_{x}$ at Euclidean distance $\delta$ from the S vertex of $S_{x}$, and $t_{x}$ the point on the NE side of $S_{x}$ at Euclidean distance $\delta$ from the N vertex. Note that $\|s_{x}-t_{x}\|_{\infty}>1/2$. Let $\mathcal{F}$} be the family of non-expansive functions $\{f_{\delta,s_{x}},f_{\delta,t_{x}}~|x\in[N]\}$.